OpenStudy (anonymous):

You have 4/6 cup of cornmeal in your cupboard. A recipe for 2/5 muffins calls for cup of cornmeal. How much cornmeal would you have left if you made the muffins?

OpenStudy (mathstudent55):

The recipe calls for 2/5 cup and you have 4/6 cup? If so, you need to subtract 4/6 - 2/5. 4/6 = 2/3 2/3 - 2/5 = You need a common denominator which is 15, (2/3)*(5/5) - (2/5)*(3/3) = 10/15 - 6/15 = 4/15 You have 4/15 cup left.

OpenStudy (anonymous):

4/15